3. If 3 books are picked at random from a shelf containing 5 novels, 3 books of poems, and a dictionary, what is the probability that
a. all books selected are novels?
b. 2 novels and 1 book of poems are selected?
1
Expert's answer
2020-10-20T18:38:35-0400

There are 9 books on the shelf. We can pick up 3 book in

"\displaystyle C(9,3) = \frac{9!}{3!6!} = \frac{7 \cdot 8 \cdot 9}{3 \cdot 2}=84" ways.

a) We can pick up 3 novels out 5 available in

"\displaystyle C(5,3) = \frac{5!}{3!2!} = \frac{5 \cdot 4}{2}=10"

So, the probability of "all books selected are novels" is

"P(A) = \frac{10}{84}= \frac{5}{42}"

b) We can pick up 2 novels and 1 book of poems in

"\displaystyle C(5,2) \cdot C(3,1) = \frac{5!}{2!3!} \cdot \frac{3!}{2!1!}= 10 \cdot 3=30"

The probability of event "2 novels and 1 book of poems are selected" is

"P(A) = \frac{30}{84}= \frac{5}{14}"


Answer: a) 5/42, b) 5/14